OpenStudy (anonymous):

X(Product) Cinnamon Roll Donut Crossiant P(X) 0.20 0.55 0.25 a) If a cinnamon roll yields $0.25 profit, a donut yields $0.20 profit, and a croissant yields $0.30 profit, how much profit will be made on each customer? b) If there are 50 customers, how much money can you expect her to make?

jimthompson5910 (jim_thompson5910):

a) is asking you to find the expected value b) is asking you to multiply what you found in part a) by 50

OpenStudy (anonymous):

could you help me find the value of part a pleas? im not sure I understand how?

jimthompson5910 (jim_thompson5910):

The expected value is the sum of all the probabilities multiplied by the values of each outcome. Basically, you multiply P(X = cinnamon roll) by V(X = cinnamon roll) So you would get 0.20 times 0.25 = 0.05 ------------------------------------------------------- You do the same for the other probabilities and profits For the donuts, you do 0.55*0.20 = 0.11 For the croissants, you do 0.25*0.30 = 0.075 ------------------------------------------------------- now you add up the three results you got above: 0.05, 0.11, 0.075 to get 0.05 + 0.11 + 0.075 = 0.235

jimthompson5910 (jim_thompson5910):

So the expected value of X is 0.235 This means they expect to make (on average) $0.235 which rounds to $0.24 per customer

jimthompson5910 (jim_thompson5910):

So $0.24 is made per customer Then you take this result and you multiply it by 50 to get the answer to part b

OpenStudy (anonymous):

Thank you so much! (:

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

So, it would be 0.24x50? which equals 12?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so she expects to make (on average) $12 for 50 customers

OpenStudy (anonymous):

Okay, it makes sense to me now! Thanks so much!

jimthompson5910 (jim_thompson5910):

sure thing, and glad it does